OpenStudy (anonymous):

Please help me with one Algebra question :) A ball is thrown into the air with an upward velocity of 28 ft/s. Its height (h) in feet after t seconds is given by the function h = –16t² + 28t + 7. How long does it take the ball to reach its maximum height? What is the ball’s maximum height? Round to the nearest hundredth, if necessary. 1.75 s; 7 ft 0.88 s; 43.75 ft 0.88 s; 17.5 ft 0.88 s; 19.25 ft

OpenStudy (anonymous):

The function is a (negative) quadratic.. Basically, it's an upside down parabola. so, if you can find the vertex, then you have found the maximum height!

OpenStudy (anonymous):

Here are 2 ways to find the vertex of a quadratic: 1) (-b/2a, f(-b/2a)) Basically, determine what -b/2a is: in this case, a=-16 b=28 and c=7 so, -b/2a = 7/8 or .88 Now, plug in .88 into the equation to determine h..

OpenStudy (anonymous):

2) find the axis of symmetry by determining where the x-intercepts are.. then, find the midpoint of those x values...

OpenStudy (anonymous):

http://www.mathplane.com/gate_5algebra_ii/quadratics_overview__notes this page contains another example within the notes...